# Mechanistic Modeling and Analysis of Transport Phenomena for Continuous Lyophilization

### v2

@@ Introduction
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, ranging from the simplest lumped capacity model [1] to the high-fidelity model considering multidimensional heat and mass transfer in detail [17]. In Ref. [25], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. Therefore, 1D modeling is preferable. Nevertheless, typical measurement techniques such as Karl Fischer titration measure the total amount of bound water, not the spatial variation. Consequently, a lumped capacity model is sometimes considered acceptable.
@@
In primary drying, the drying rate is mainly governed by the rate of sublimation, which makes the modeling typically more straightforward than for secondary drying. In secondary drying, the two important mechanisms are (1) vapor phase transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Some previous models for secondary drying consider both phenomena [14, 15, 16, 17], whereas the more recently published models tend to consider only the desorption part [23, 4, 24, 25, 36]. Simulation results in these studies show that the model prediction is highly accurate even when the vapor phase transport is omitted, suggesting that desorption is the actual limiting step. In Ref. [37], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying. Evidence from both simulation and experiment suggest that only the desorption part is necessary.
@@
Optimal modeling strategies
@@
Our goal is to select modeling strategies that can capture critical phenomena in the process and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Fig. 2. The freezing model relies on the lumped capacity method that considers nucleation and supercooling. The primary drying model simulates heat transfer in 1D (vertical direction) with mass transfer incorporated as a boundary condition at the sublimation front. Finally, the secondary drying model captures heat transfer in 1D (vertical direction) and desorption of bound water. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
@@
Model for freezing
For the design proposed by Ref. [7], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [38]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf.
Our model (Fig. 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
Preconditioning
@@
where $T(t)$ is the temperature, $t$ is time, $m$ is the total mass, $C_{p}$ is the specific heat capacity (per mass), $Q$ is the total heat transfer between the vial surface and environment, the subscripts ‘w’ and ‘s’ refer to the water (solvent) and solid (solute) phases, and the subscripts ‘s1’, ‘s2’ and ‘s3’ denote the top, bottom, and side of the vial, respectively. The initial conditions for (1) are
@@
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, $\mathcal{F}$ is the transfer factor (see the definition in Section 3.5), and the subscripts ‘u’, ‘b’, ‘c’, and ‘g’ denote the upper surface of the chamber, bottom shelf, chamber walls, and cold gas, respectively. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. In lyophilization, the bottom shelf and gas temperatures ($T_{\textrm{b}}$ and $T_{\textrm{g}}$) typically vary with time and are specified by the freezing/drying protocol. The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [39]. In this article, we assume that both temperatures are approximately constant as reported by Refs. [40, 19]. In addition, $T_{\textrm{u}}\approx T_{\textrm{c}}$ due to no additional heat source/sink at these locations. Nevertheless, our model is not restricted to these assumptions; i.e., temperature-dependent data can be fed to the model (if available).
Next, (4)–(6) are simplified further to match the experimental setup in Ref. [7], whose data are used for our model validation. In Ref. [7], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection. In this setup, the upper surface and local gas in that area are not cooled, so the gas temperature is assumed to be equal to the surface temperature $T_{\textrm{u}}$. Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of Ref. [31] and Section 3.5). Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. As a result, $Q_{\textrm{s1}}$ and $Q_{\textrm{s2}}$ are
@@
At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for (6).
Note that the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by Ref. [7] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying (4)–(6) even further.
Vacuum-induced surface freezing
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until its completion at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure to evaporate a small amount of liquid solution from the product, abruptly decreasing the temperature and promoting nucleation. A model for VISF is not readily available in the literature but can be derived from the mechanistic understanding of simultaneous heat and mass transfer during evaporation, e.g., as described in Ref. [31]. First, consider the mass transfer part. The evaporation rate of water at the surface for a nonvolatile solute is
@@
where $h_{m}$ is the mass transfer coefficient, $x_{\textrm{w,sat}}$ is the mole fraction of water at the liquid-vapor interface (equilibrium), and $x_{\textrm{w,c}}$ is the mole fraction of water in the chamber (environment). The subscripts ‘sat’ and ‘c’ here denote the equilibrium condition and environment, respectively. To calculate the mass fraction of water for (9), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’. By assuming the ideal gas law,
@@
where $p_{\textrm{t}}$ is the total pressure, $p_{\textrm{w}}$ is the partial pressure of water, and $M$ is the molar mass. If the environment contains only nitrogen or inert gas, $p_{\textrm{w,c}}=0$. The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [41],
@@
Note that, if the VISF process is done properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
Next, consider the heat transfer part. The energy balance equation can be modified from (1) as
@@
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [41]
@@
In (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
The initial conditions for (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by Ref. [12]. In this work, we use the term nucleation to denote the first nucleation where the temperature of supercooled liquid almost instantaneously increases to the equilibrium/freezing point, which is caused by the heat released from the fraction of liquid being frozen/solidified. On the other hand, we use the term solidification to denote the phase transition from liquid to solid (ice formation) after that first nucleation.
@@
where $T_{\textrm{f,w}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving both (15) and (16) simultaneously.
Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes. Since the nucleation process is nearly instantaneous as explained by Ref. [12], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
@@
The probability defined by (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. Subsequently, follow the exact same procedure described by (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
Solidification
After the first nucleation completes, the solidification (aka ice formation) step starts at $t_{\textrm{f3}}$. For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [12]; we also discuss this in the context of the Biot number in Appendix A. The lumped capacity model is sufficiently accurate and highly computationally efficient, but it does not represent the correct physics of ice formation. By treating the entire solid/liquid as a lumped object during solidification, it implies that homogeneous nucleation is assumed. Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [42]. Consequently, the solidification process generally starts from the vial surfaces, i.e., the outer and bottom surfaces in this case. To take this fact into account, a rigorous 2D heat transfer model could be considered [13]. In such cases, the physics is incorporated more accurately, but the major drawback of 2D modeling is its high computational cost and numerical complexity.
Our strategy here is to combine the advantages of the lumped capacity and 2D models together, resulting in a model that is computationally efficient, provides accurate prediction, and represents the correct physics; we denote it as a hybrid lumped capacity model. In this case, ice formation is assumed to initiate from the side and bottom surfaces (similar to a 2D model), in which the ice layer is treated as an additional heat transfer resistance between the cold gas and unfrozen water. Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [31]. Physically, this is equivalent to assuming that all the heat input is used for phase transition only. The unfrozen water and solute are treated as a lumped object as its temperature follows the equilibrium temperature. A schematic diagram showing our model for the solidification step is shown in Fig. 4
@@
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in (7), (8), and (6), respectively. During the solidification step, the amount of water $m_{\textrm{w}}(t)$ and ice $m_{\textrm{i}}(t)$ varies with time, which also affects the volume and surface area of the product. The relation between $m_{\textrm{w}}$ and $m_{\textrm{i}}$ is
@@
The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in (8) and (6) can be replaced by the overall heat transfer coefficients
@@
where $r_{\textrm{o}}$, $r$ and $l$ are as defined in Fig. 4 and $k_{\textrm{i}}$ is the thermal conductivity of ice. For the radiation part, the linearization technique described in Section 3.2.1 can be applied, and so the equation is written in the form of Newton’s law of cooling. Note that the overall heat transfer coefficients in (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$
The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely (17), (19), and (21).
Finally, define $t_{\textrm{f4}}$ as the time when solidification completes. The criterion for complete solidification is defined as
@@
where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [12].
Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in Ref [12]) while describing the physics in a more accurate way as in the 2D model (e.g., in Ref [13]). Final simulations results from these models are almost identical for lyophilization of unit doses, and so there is no concern in terms of model accuracy.
Cooling
The final step is to ensure that the product temperature is at the desire value before starting the drying process. This cooling step starts at $t_{\textrm{f4}}$ and completes at $t_{\textrm{f5}}$. The model for the final cooling step is relatively simple, with the only energy balance equation

### v3

@@ 3.1.3 Modeling strategies for the secondary drying step
Secondary drying is similar to primary drying, with the liquid mainly removed via desorption instead of sublimation. Several modeling strategies are available, ranging from the simplest lumped capacity model [Fissore2018Review] to the high-fidelity model considering multidimensional heat and mass transfer in detail [Sheehan1998Modeling]. In [Yoon2021Sec0D1D3D], it was shown that the accuracy of 1D modeling is comparable to that of multidimensional modeling while being much less complicated, whereas a lumped capacity model could produce significant errors. Therefore, 1D modeling is preferable. Nevertheless, typical measurement techniques such as Karl Fischer titration measure the total amount of bound water, not the spatial variation. Consequently, a lumped capacity model is sometimes considered acceptable.
@@
In primary drying, the drying rate is mainly governed by the rate of sublimation, which makes the modeling typically more straightforward than for secondary drying. In secondary drying, the two important mechanisms are (1) vapor phase transport through the porous dried product and (2) desorption of bound water from the solid surface of the dried product. Some previous models for secondary drying consider both phenomena [Litchfield1979Model, Liapis1994Original, Sadikoglu1997Modeling, Sheehan1998Modeling], whereas the more recently published models tend to consider only the desorption part [Fissore2011SecDryingMonitor, Fissore2015Review, Sahni2017Simplified, Yoon2021Sec0D1D3D, Srisuma2024Obs]. Simulation results in these studies show that the model prediction is highly accurate even when the vapor phase transport is omitted, suggesting that desorption is the actual limiting step. In [Pikal1990SecDrying_Kinetics], a detailed experimental study was conducted, concluding that desorption is the rate-limiting step for mass transfer in secondary drying. Evidence from both simulation and experiment suggest that only the desorption part is necessary.
@@
3.1.4 Optimal modeling strategies
@@
Our goal is to select modeling strategies that can capture the critical phenomena in lyophilization and produce sufficiently accurate results while keeping the model complexity, computational cost, and number of parameters at a minimum. The resulting models should be sufficiently accurate and efficient to be used for general process design (e.g., input/output design, heat/material balance, selecting operating conditions), process optimization, state estimation, and real-time model-based control (e.g., model predictive control). For detailed process and equipment design, a high-fidelity model is needed, which is beyond the scope of this work.
From the discussion in Sections 3.1.13.1.3 and Appendix A, the selected modeling strategies are highlighted in Figure 2. The freezing model relies on the lumped capacity method that considers nucleation and supercooling. The primary drying model simulates heat transfer in 1D (vertical direction) with mass transfer incorporated as a boundary condition at the sublimation front. Finally, the secondary drying model captures heat transfer in 1D (vertical direction) and desorption of bound water. The models for continuous lyophilization developed in the subsequent sections are based on these strategies, with proper modifications for the suspended-vial configuration. Sections 3.2, 3.3, and 3.4 describe the models for freezing, primary drying, and secondary drying, respectively. Additionally, to supplement the mechanistic understanding of the proposed models, Section 3.5 specifically discusses modeling strategies and theories underpinning convection and thermal radiation, the key heat transfer mechanisms in suspended vials.
@@
3.2 Model for freezing
For the design proposed by [Capozzi2019ContLyo_SuspendedVials], suspended vials are cooled by using the cryogenic gas at a controlled temperature and flow rate. Besides, a dedicated chamber is added to control the nucleation temperature using the vacuum-induced surface freezing (VISF) technology [Kramer2002VISF]. With this setup, all surfaces of a vial experience a similar heat transfer condition, which is different from batch lyophilization where only the bottom surface is cooled by the cooling shelf.
Our model (Figure 3A) for freezing of suspended vials is divided into five main steps: (1) preconditioning, (2) VISF, (3) nucleation, (4) solidification or ice formation, and (5) final cooling.
3.2.1 Preconditioning
@@
where $T(t)$ is the temperature, $t$ is time, $m$ is the total mass, $C_{p}$ is the specific heat capacity (per mass), $Q$ is the total heat transfer between the vial surface and environment, the subscripts ‘w’ and ‘s’ refer to the water (solvent) and solid (solute) phases, and the subscripts ‘s1’, ‘s2’ and ‘s3’ denote the top, bottom, and side of the vial, respectively. The initial conditions for Equation (1) are
@@
where $h$ is the heat transfer coefficient, $A_{z}$ is the cross sectional area of the product, $A_{r}$ is the side surface area of the product, $\sigma$ is the Stefan-Boltzmann constant, $\mathcal{F}$ is the transfer factor (see the definition in Section 3.5), and the subscripts ‘u’, ‘b’, ‘c’, and ‘g’ denote the upper surface of the chamber, bottom shelf, chamber walls, and cold gas, respectively. In this work, the vial/product is modeled as a cylinder of diameter $d$, and hence the cross sectional and surface area can be calculated from the given volume and diameter. In lyophilization, the bottom shelf and gas temperatures ($T_{\textrm{b}}$ and $T_{\textrm{g}}$) typically vary with time and are specified by the freezing/drying protocol. The upper surface and wall temperatures ($T_{\textrm{u}}$ and $T_{\textrm{c}}$), however, are not usually measured and thus can be estimated from data [Srisuma2024Rad]. In this article, we assume that both temperatures are approximately constant as reported by [Gan2005WallTemp, Veraldi2008SimBatchModels]. In addition, $T_{\textrm{u}}\approx T_{\textrm{c}}$ due to no additional heat source/sink at these locations. Nevertheless, our model is not restricted to these assumptions; i.e., temperature-dependent data can be fed to the model (if available).
Next, Equations (4)–(6) are simplified further to match the experimental setup in [Capozzi2019ContLyo_SuspendedVials], whose data are used for our model validation. In [Capozzi2019ContLyo_SuspendedVials], the gas inside the chamber was cooled by the bottom shelf, and this cold gas was subsequently used to cool the vials, hence natural convection. In this setup, the upper surface and local gas in that area are not cooled, so the gas temperature is assumed to be equal to the surface temperature $T_{\textrm{u}}$. Consequently, the two terms can be combined and written in the form of Newton’s law of cooling, with $h_{\textrm{s1}}$ combining the effects of both thermal radiation and natural convection, which can be done by linearizing the fourth-order term of the radiation part (see more details in Section 1.3.2 of [Mills1995HeatTransfer] and Section 3.5). Similarly, the gas at the bottom surface is cooled by the bottom shelf, so we assume that both gas and bottom shelf have the same temperature $T_{\textrm{g}}$, with $h_{\textrm{s2}}$ combining both radiation and convection. As a result, $Q_{\textrm{s1}}$ and $Q_{\textrm{s2}}$ are
@@
At the side surfaces, there are natural convection from the cold gas at $T_{\textrm{g}}$ and thermal radiation from the chamber wall $T_{\textrm{c}}$, and so no simplification is needed for Equation (6).
Note that the expressions for $Q_{\textrm{s1}},Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ are dependent on the system. Another design proposed by [Capozzi2019ContLyo_SuspendedVials] is to flush the cryogenic gas directly into the chamber. In that case, the dominant heat transfer mode is forced convection, which is much stronger than thermal radiation in this low temperature region. Thus, the effect of thermal radiation might be omitted, simplifying Equations (4)–(6) even further.
3.2.2 Vacuum-induced surface freezing
At the end of the preconditioning step, VISF is initiated at $t_{\textrm{f1}}$ and proceeds until its completion at $t_{\textrm{f2}}$. The key idea of VISF is to reduce the total pressure to evaporate a small amount of liquid solution from the product, abruptly decreasing the temperature and promoting nucleation. An example of available models for VISF can be found in [Pisano2017VISF], which relies on the concept of the condensing/evaporating efficiency. A similar approach is used in this work, but the model is derived from the fundamental of heat and mass transfer during evaporation, e.g., as described in [Mills1995HeatTransfer], and thus the associated parameters are the heat and mass transfer coefficients instead of the condensing/evaporating efficiency. In addition, the stochastic nature of ice nucleation is also incorporated into the VISF model (described later in Section 3.2.3), allowing for the comparison between spontaneous nucleation and VISF. First, consider the mass transfer part. The evaporation rate of water at the surface for a nonvolatile solute is
@@ 3.2.2 Vacuum-induced surface freezing
where $h_{m}$ is the mass transfer coefficient, $x_{\textrm{w,sat}}$ is the mole fraction of water at the liquid-vapor interface (equilibrium), and $x_{\textrm{w,c}}$ is the mole fraction of water in the chamber (environment). The subscripts ‘sat’ and ‘c’ here denote the equilibrium condition and environment, respectively. To calculate the mass fraction of water for Equation (9), we first assume that there are two gas/vapor components during VISF, namely (1) water and (2) nitrogen or inert gas, denoted by the subscript ‘in’. By assuming the ideal gas law,
@@
where $p_{\textrm{t}}$ is the total pressure, $p_{\textrm{w}}$ is the partial pressure of water, and $M$ is the molar mass. If the environment contains only nitrogen or inert gas, $p_{\textrm{w,c}}=0$. The saturation pressure $p_{\textrm{w,sat}}$ is a function of temperature, that is [Smith2018Thermo],
@@
Note that, if the VISF process is carried out properly, the amount of water that vaporizes is usually very small and thus does not significantly impact the overall thermophysical properties.
Next, consider the heat transfer part. The energy balance equation can be modified from Equation (1) as
@@
where the last term on the right-hand side is the amount of heat removed via evaporation and $\Delta H_{\textrm{vap}}$ is the heat of vaporization, which can be approximated by [Smith2018Thermo]
@@
In Equation (13), $m_{\textrm{w}}(t)$ is time-dependent due to evaporation, which also means that the side surface area $A_{r}$ changes with time.
The initial conditions for Equations (9) and (13) are the final mass of water and temperature at the end of the preconditioning step.
3.2.3 Nucleation
The nucleation step starts at the end of VISF ($t_{\textrm{f2}}$) for controlled nucleation or at the end of the preconditioning step ($t_{\textrm{f1}}$) for uncontrolled nucleation, and then completes at $t_{\textrm{f3}}$. Our modeling strategy for the nucleation step described below is based on the state-of-the-art freezing model proposed by [Deck2022FreezingLumped]. In this work, we use the term nucleation to denote the first nucleation where the temperature of supercooled liquid almost instantaneously increases to the equilibrium/freezing point, which is caused by the heat released from the fraction of liquid being frozen/solidified. On the other hand, we use the term solidification to denote the phase transition from liquid to solid (ice formation) after that first nucleation.
@@ 3.2.3 Nucleation
where $T_{\textrm{f,w}}$ is the freezing point of pure water, $K_{\textrm{f}}$ is the molal freezing-point depression constant, and $M_{\textrm{s}}$ is the molar mass of a solute. The two unknowns $T_{\textrm{f,l}}$ and $m_{\textrm{i,n}}$ can be obtained by solving Equations (15) and (16) simultaneously.
Finally, define $t_{\textrm{f3}}$ as the time when the nucleation process completes. Since the nucleation process is nearly instantaneous as explained by [Deck2022FreezingLumped], $t_{\textrm{f3}}$ is set to $t_{\textrm{f2}}$, with the relations
@@
The probability defined by Equation (24) is calculated cumulatively at every time step $\Delta t$ until the first nucleation occurs, which is marked as the end of the preconditioning step $t_{\textrm{f1}}$. Subsequently, follow the exact same procedure described by Equations (15)–(21), with $t_{\textrm{f1}}$ replacing $t_{\textrm{f2}}$.
3.2.4 Solidification
After the first nucleation completes, the solidification (aka ice formation) step starts at $t_{\textrm{f3}}$. For solidification, the lumped capacity model was shown to provide accurate prediction of the temperature and freezing time for typical vial sizes [Deck2022FreezingLumped]; we also discuss this in the context of the Biot number in Appendix A. The lumped capacity model is sufficiently accurate and highly computationally efficient, but it does not represent the correct physics of ice formation. By treating the entire solid/liquid as a lumped object during solidification, it implies that homogeneous nucleation is assumed. Nevertheless, it is commonly known that heterogeneous nucleation is a more frequently observed phenomenon during ice formation [Glatz2018Nucleation]. Consequently, the solidification process generally starts from the vial surfaces, i.e., the outer and bottom surfaces in this case. To take this fact into account, a rigorous 2D heat transfer model could be considered [Deck2024Freezing2D]. In such cases, the physics is incorporated more accurately, but the major drawback of 2D modeling is its high computational cost and numerical complexity.
Our strategy here is to combine the advantages of the lumped capacity and 2D models together, resulting in a model that is computationally efficient, provides accurate prediction, and represents the correct physics; we denote it as a hybrid lumped capacity model. In this case, ice formation is assumed to initiate from the side and bottom surfaces (similar to a 2D model), in which the ice layer is treated as an additional heat transfer resistance between the cold gas and unfrozen water. Additionally, heat conduction in the ice layer is assumed to be quasi steady, which is a reasonable assumption for problems associated with phase change in general as the sensible heat is much smaller than the latent heat [Mills1995HeatTransfer]. Physically, this is equivalent to assuming that all the heat input is used for phase transition only. The unfrozen water and solute are treated as a lumped object as its temperature follows the equilibrium temperature. A schematic diagram showing our model for the solidification step is shown in Figure 4
@@ 3.2.4 Solidification
with $Q_{\textrm{s1}}$, $Q_{\textrm{s2}}$, and $Q_{\textrm{s3}}$ as defined in Equations (7), (8), and (6), respectively. During the solidification step, the amount of water $m_{\textrm{w}}(t)$ and ice $m_{\textrm{i}}(t)$ varies with time, which also affects the volume and surface area of the product. The relation between $m_{\textrm{w}}$ and $m_{\textrm{i}}$ is
@@
The ice layer becomes an additional heat transfer resistance, and so the heat transfer coefficients $h_{\textrm{s2}}$ and $h_{\textrm{s3}}$ in Equations (8) and (6) can be replaced by the overall heat transfer coefficients
@@
where $_{\textrm{o}}, rr$ and $l$ are as defined in Figure 4 and $k_{\textrm{i}}$ is the thermal conductivity of ice. For the radiation part, the linearization technique described in Section 3.2.1 can be applied, and so the equation is written in the form of Newton’s law of cooling. Note that the overall heat transfer coefficients in Equations (28) and (29) are defined based on the outer surface area to be consistent with the original expressions of $Q_{\textrm{s2}}$ and $Q_{\textrm{s3}}$ .
The initial conditions for $m_{\textrm{w}}$, $m_{\textrm{i}}$, and $T$ are the final conditions of the nucleation step, namely Equations (17), (19), and (21). Finally, define $t_{\textrm{f4}}$ as the time when solidification completes. The criterion for complete solidification is defined as
@@
where the coefficient could vary between 0.85 and 0.95 without significantly changing the final results [Deck2022FreezingLumped].
Our hybrid lumped capacity model presented in this section has the same model complexity and computational cost as those of the original lumped capacity model (e.g., in [Deck2022FreezingLumped]) while describing the physics in a more accurate way as in the 2D model (e.g., in [Deck2024Freezing2D]). Final simulations results from these models are almost identical for lyophilization of unit doses, and so there is no concern in terms of model accuracy.
3.2.5 Cooling
The final step is to ensure that the product temperature is at the desire value before starting the drying process. This cooling step starts at $t_{\textrm{f4}}$ and completes at $t_{\textrm{f5}}$. The energy balance is